OpenStudy (anonymous):

Find the real number solutions of this equation and show the work!!! x^4-32=14x^2 Find the real number solutions of this equation and show the work!!! x^4-32=14x^2 @Mathematics

OpenStudy (anonymous):

Let x^2 be a x^4 - 32 = 14x^2 x^4 - 14x^2 - 32 = 0 Then you apply the x^2 = a part a^2 - 14a - 32 = 0 Factoring (a-16)(a+2) Then you put a back into x^2 (x^2 - 16)(x^2 + 2) x = 4, -square root 2

OpenStudy (anonymous):

okay, but after factoring (x^2-16)(x^2+2) dont you get: x= -1, +/- 4?

OpenStudy (anonymous):

oh yea, the first part, it's 4 and -4. For (x^2 + 2), you're trying to make that into 0 as well. (-1)^2 + 2 would be 3, and the square will turn it into a positive no matter what, so.. I think it's 0,+/- 4

OpenStudy (anonymous):

wait it's not 0.. Not sure about x^2 + 2.. Havent done these in ages x.x